### v1

# Numerical valuation of European options under two-asset infinite-activity exponential Lévy models

Abstract
We propose a numerical method for the valuation of European-style options under two-asset infinite-activity exponential Lévy models. Our method extends the effective approach developed by Wang et al. (2007) for the 1-dimensional case to the 2-dimensional setting and is applicable for general Lévy measures under mild assumptions. A tailored discretization of the non-local integral term is developed, which can be efficiently evaluated by means of the fast Fourier transform. For the temporal discretization, the semi-Lagrangian $\theta$-method is employed in a convenient splitting fashion, where the diffusion term is treated implicitly and the integral term is handled explicitly by a fixed-point iteration. Numerical experiments for put-on-the-average options under Normal Tempered Stable dynamics reveal favourable second-order convergence of our method whenever the exponential Lévy process has finite-variation.
1 Introduction
The accurate valuation of derivative securities in modern financial markets requires modeling techniques capable of capturing empirical irregularities in asset price dynamics. Classical models based on Brownian motion, such as the Black–Scholes model, rely on continuous-path diffusion and fail to reflect important stylized facts, such as heavy tails and skewness in log-returns. This has motivated the use of Lévy processes in the last decades, which naturally offers a richer class of models for asset dynamics. Among various Lévy models, the Normal Inverse Gaussian (NIG) process has emerged as a parsimonious and effective choice to capture such characteristics. Among others, Rydberg (1997) shows how the NIG model provides a significantly better statistical fit to equity return data compared to classical Gaussian-based models. Lévy models allow for a more realistic representation of market risk and are therefore natural candidates for use in option pricing models.
In this paper, we propose a numerical method for pricing European-style financial derivatives written on two underlying assets, whose dynamics are driven by a 2-dimensional Lévy process, with particular focus on infinite activity processes. Financial pricing under jump-diffusion models can be approached through various methodologies, such as Monte Carlo simulation, Fourier-based methods, and partial integro-differential equations (PIDEs). Monte Carlo methods are flexible and easy to implement, but they suffer from slow convergence. Fourier-based methods, such as in Jackson et al. (2008) and Ruijter & Oosterlee (2012), can be applied when the characteristic exponent of the process is known, and they can achieve exponential convergence. Numerical methods for PIDEs, such as in Cont & Voltchkova (2005), d’Halluin et al. (2005), Wang et al. (2007), Clift & Forsyth (2008), Salmi et al. (2014), Kaushansky et al. (2018), Boen & in ’t Hout (2021) and in ’t Hout & Lamotte (2023), can instead be applied when the Lévy measure is known, and do not require knowledge of the characteristic exponent. They are applicable to a wide variety of financial derivatives.
The numerical method derived in this paper focuses on the case where the underlying 2-dimensional Lévy process exhibits infinite-activity, meaning that an infinite number of jumps occur over any finite time horizon. In this setting, particular care must be taken in the discretization of the non-local 2-dimensional integral term in the PIDE near the origin, where the Lévy measure becomes singular. Indeed, classical quadrature formulae fail to yield the desired second-order convergence.
The main contribution of this paper is an extension of the effective numerical solution approach of Wang et al. (2007) from the 1-dimensional to the 2-dimensional setting. Here a key idea, originally introduced in Asmussen & Rosiński (2001) and Cont & Voltchkova (2005), is to replace the small jumps with an artificial diffusion term. This substitution enables the development of a tailored quadrature scheme, which restores the desired order of convergence of the entire numerical scheme. For the efficient evaluation of the discretized integral operator, a fast Fourier transform (FFT) algorithm is constructed. For the temporal discretization, the semi-Lagrangian $\theta$-method is considered. Here operator splitting is applied, where the diffusion term is treated implicitly and the integral term is handled explicitly by a fixed-point iteration. For the large linear system in each time step, the BiCGSTAB iterative solver is used.
An outline of this paper is as follows. In Section 2, we introduce the market model and the PIDE for the derivative pricing. In Section 3 the proposed numerical scheme is derived. Numerical experiments are discussed in Section 4. The final Section 5 contains our conclusions.
2 Model framework
2.1 Market model
Let $\left(\Omega,\mathcal{F},\left(\mathcal{F}_{t}\right)_{t\in\left[0,T\right]},\mathbb{P}\right)$ be a filtered probability space, for some given $T>0$. We consider an arbitrage free market characterized by a constant (instantaneous) risk-free interest rate $r$ and an equivalent martingale measure $\mathbb{Q}\sim\mathbb{P}$. We assume that there exist two risky assets, whose prices are modeled by the 2-dimensional stochastic process $X=\left(X^{\left(1\right)},X^{\left(2\right)}\right)$ that solves the following stochastic differential equation
$dX\left(t\right)=\mu\left(X\left(t\right)\right)dt+\Sigma\left(X\left(t\right)\right)dW\left(t\right)+\int_{\mathbb{R}_{*}^{2}}\gamma\left(z,X\left(t_{-}\right)\right)\tilde{\Pi}\left(dt,dz\right)\qquad\left(t\in\left(0,T\right]\right)$ (2.1)
for some non-negative initial value $X\left(0\right)$. In (2.1), $W$ denotes a standard 2-dimensional Wiener process and $\tilde{\Pi}$ is a compensated Poisson random measure with Lévy measure $\ell$ over $\mathbb{R}_{*}^{2}=\mathbb{R}^{2}\setminus\left\{0\right\}$. Both are directly defined under $\mathbb{Q}$ and are mutually independent.
The functions $\mu:\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2}$, $\Sigma:\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2\times 2}$ and $\gamma:\mathbb{R}^{2}\times\mathbb{R}_{\geq 0}^{2}\rightarrow\mathbb{R}^{2}$ are called drift, diffusion and jump function (or term) respectively, where $\mathbb{R}_{\geq 0}^{2}=\left\{x\in\mathbb{R}^{2}:x^{\left(i\right)}\geq 0\text{ for }i=1,2\right\}$. In this paper, we consider the case of the well-known exponential Lévy process, i.e. where the coordinates of the coefficient functions are defined as follows
$\mu^{\left(i\right)}\left(x\right)$ $=x^{\left(i\right)}r$ (2.2)
$\left(\Sigma\Sigma^{\top}\right)^{\left(i,j\right)}\left(x\right)$ $=x^{\left(i\right)}x^{\left(j\right)}\left(\sigma\sigma^{\top}\right)^{\left(i,j\right)}$ (2.3)
$\gamma^{\left(i\right)}\left(z,x\right)$ $=x^{\left(i\right)}\left(e^{z^{\left(i\right)}}-1\right),$ (2.4)
where $\sigma\sigma^{\top}$ is a constant positive definite symmetric $2\times 2$ matrix and $\sigma$ denotes the volatility matrix. Here, $\Sigma\Sigma^{\top}\left(x\right)$ is a shorthand notation for the matrix product $\Sigma\left(x\right)\Sigma^{\top}\left(x\right)$.
Let $\left\|\cdot\right\|$ be any given norm on $\mathbb{R}^{2}$. In this paper we assume that $\ell$ is absolutely continuous, has finite variance, i.e.
$\int_{\mathbb{R}}z^{2}\ell\left(dz\right)<\infty$
and there exist constants $A_{\ell}<2$, $B_{\ell}>2$ such that for any given $h>0$ there is $C_{\ell}\left(h\right)>0$ such that
$\begin{cases}\ell\left(z\right)\leq C_{\ell}\left(h\right)\left\|z\right\|^{-A_{\ell}-2}&\text{for any }z\text{ such that }\left\|z\right\|\in\left(0,h\right]\\ \ell\left(z\right)=O\left(e^{-B_{\ell}\left\|z\right\|}\right)&\text{as }\left\|z\right\|\rightarrow\infty.\end{cases}$ (2.5)
The number $A_{\ell}$ governs the activity and variation of the process: $X$ is of finite-activity if $A_{\ell}<0$, since $\int_{\mathbb{R}^{2}}\ell\left(dz\right)<\infty$; it is of finite-variation if $A_{\ell}<1$, since $\int_{\left\|z\right\|<\epsilon}\left\|z\right\|\ell\left(dz\right)<\infty$ for any $\epsilon>0$. The number $B_{\ell}$ characterizes the exponential decay of $\ell$ at infinity. Since the process $X$ has finite moments of all orders up to $k\in\mathbb{N}$ if and only if $\int_{\left\|z\right\|>\epsilon}e^{k\left\|z\right\|}\ell\left(dz\right)<\infty$ for any $\epsilon>0$, then $k<B_{\ell}$ provides a necessary condition of it. Following Applebaum (2004, Chapter 6), the stronger condition $B_{\ell}\geq 2$ is necessary to guarantee the existence of a unique solution with finite variance to the stochastic differential equation (2.1). Most of the common Lévy processes in finance satisfy the conditions (2.5), such as the Kou, Carr-Geman-Madan-Yor (CGMY), Variance Gamma (VG) and Normal Inverse Gaussian (NIG) models.
In this work, we focus on the case of 2-dimensional Normal Tempered Stable (NTS) processes. These are obtained by subordinating a 2-dimensional arithmetic Brownian motion with a Tempered Stable subordinator. A detailed construction of the NTS process together with its main properties is provided in Appendix A. The choice of this class of processes is motivated by two reasons. First, bivariate VG and NIG processes arise as particular cases. Second, the associated Lévy measure satisfies the conditions (2.5) with constant $A_{\ell}=2\alpha$, where $\alpha$ is the key model parameter. The NTS framework provides a convenient and flexible setting for the purposes of this paper.
2.2 Initial boundary value problem for derivatives pricing
By the fundamental theorem of asset-pricing, the value at time $t\in\left[0,T\right]$ of an European-style [^1] Means a financial derivative with no intermediate cash flows. financial derivative of $X$ with maturity $T$ is represented by the stochastic process $P$ such that
$P\left(t\right)=\mathbb{E}^{\mathbb{Q}}\left[\phi\left(X\left(T\right)\right)e^{-r\left(T-t\right)}\mid\mathcal{F}_{t}\right]$
where $\phi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ denotes the pay-off function and $\mathbb{E}^{\mathbb{Q}}\left[\cdot\mid\mathcal{F}_{t}\right]$ is the $\mathcal{F}_{t}$-conditional expected value (i.e. knowing the history of the asset prices up to $t$) under $\mathbb{Q}$.
Let $\mathcal{A}$ be the infinitesimal generator of $X$ (see Applebaum (2004), Garroni & Menaldi (1992) and Øksendal & Sulem (2019)), defined in matrix notation as [^2] By expanding the term, we obtain the common notation used for $\mathcal{A}$, which is $\mathcal{A}u\left(x,t\right)=$ $\sum_{i=1}^{2}\mu^{\left(i\right)}\left(x\right)\frac{\partial u}{\partial x^{\left(i\right)}}\left(x,t\right)+\frac{1}{2}\sum_{i,j=1}^{2}\left(\Sigma\Sigma^{\top}\right)^{\left(i,j\right)}\left(x\right)\frac{\partial^{2}u}{\partial x^{\left(i\right)}\partial x^{\left(j\right)}}\left(x,t\right)$ $+\int_{\mathbb{R}_{*}^{2}}\left(u\left(x+\gamma\left(z,x\right),t\right)-u\left(x,t\right)-\sum_{i=1}^{2}\gamma^{\left(i\right)}\left(z,x\right)\frac{\partial u}{\partial x^{\left(i\right)}}\left(x,t\right)\right)\ell\left(dz\right).$
$\mathcal{A}u\left(x,t\right)=\mu\left(x\right)^{\top}u_{x}\left(x,t\right)+\frac{1}{2}\mathbf{1}^{\top}\left(u_{xx}\left(x,t\right)\circ\Sigma\Sigma^{\top}\left(x\right)\right)\mathbf{1}+\int_{\mathbb{R}_{*}^{2}}f\left(z,x,t\right)\ell\left(dz\right)$ (2.6)
where $\mathbf{1}=\left[1,1\right]^{\top}$, the symbol $\circ$ denotes the Hadamard (element-wise) product [^3] In this paper, we use the convention $AB\circ CD=\left(AB\right)\circ\left(CD\right)$, for any suitable matrices $A,B,C,D$. and
$f\left(z,x,t\right)=u\left(x+\gamma\left(z,x\right),t\right)-u\left(x,t\right)-\gamma\left(z,x\right)^{\top}u_{x}\left(x,t\right).$ (2.7)
If there exists a function $u:\mathbb{R}_{\geq 0}^{2}\times\left[0,T\right]\rightarrow\mathbb{R}$ that solves the following initial value problem for a partial integro-differential equation (PIDE)
$\begin{cases}u_{t}\left(x,t\right)=\mathcal{A}u\left(x,t\right)-ru\left(x,t\right)&\text{for any }\left(x,t\right)\in\mathbb{R}_{\geq 0}^{2}\times\left(0,T\right]\\ u\left(x,0\right)=\phi\left(x\right)\end{cases}$ (2.8)
then the value of the financial derivative is given by
$P\left(t\right)=u\left(X\left(t\right),T-t\right).$
Note that $u$ also satisfies the PIDE at the boundary of the $x$-domain, as in the case of option pricing with the Black–Scholes model.
3 Numerical scheme
In this section, we derive the numerical scheme proposed for problem (2.8).
The method consists of three main steps: integral discretization, spatial discretization, and temporal discretization. By discretization, we mean that the pertinent integro/differential operators are approximated on a given finite set of grid points. The adjectives indicate the variable being discretized: integral for $z$, spatial for $x$, and temporal for $t$.
The integral discretization yields an approximation to the integral term in (2.8) for any given pair $\left(x,t\right)\in\mathbb{R}_{\geq 0}^{2}\times\left[0,T\right]$. The quadrature formula that we derive is inspired by the ideas in Wang et al. (2007) and Cont & Voltchkova (2005), where the singular part of the integral near the origin $z=0$ is approximated by a diffusion (second-order spatial derivative). The integral discretization leads to the approximate problem (3.8) where the integral in (2.8) has been replaced by a summation term.
The spatial discretization concerns the diffusion and summation terms in (3.8). For the diffusion term, a standard second-order central finite difference scheme is applied on a suitable nonuniform spatial grid. For the summation term, a direct valuation on the spatial grid is computationally too expensive. For the efficient treatment of this term, we shall extend the FFT-based approach by Wang et al. (2007).
The temporal discretization is done by the semi-Lagrangian $\theta$-method. The semi-Lagrangian approach is chosen to take into account that problem (3.8) can be convection-dominated. In each time step of the semi-Lagrangian $\theta$-method, the summation term appears in an implicit manner. To effectively handle this term, a fixed-point iteration is employed.
3.1 Integral discretization
When the Lévy measure is singular, it is not possible to apply classical quadrature formulae, such as the midpoint or the trapezoidal rule. In fact, in this case the error will blow up as the number of quadrature points increases. To address this problem, we develop in this subsection a different quadrature formula.
First, it is useful to investigate the behaviour of $f$, defined in (2.7), around the origin with respect to $z$. For any given point $\left(x,t\right)\in\mathbb{R}_{\geq 0}^{2}\times\left[0,T\right]$, the Taylor approximation of the function $z\mapsto f\left(z,x,t\right)$ at $z=0$ is given by
$f\left(z,x,t\right)=f\left(0,x,t\right)+z^{\top}f_{z}\left(0,x,t\right)+\frac{1}{2}z^{\top}f_{zz}\left(0,x,t\right)z+\varepsilon\left(z,x,t\right)\qquad\text{as }\left\|z\right\|\rightarrow 0^{+},$
where $f_{z}$ and $f_{zz}$ are the gradient and the Hessian of $f$ with respect to $z$. Here, $\varepsilon$ denotes the remainder and is such that $\varepsilon\left(z,x,t\right)=O\left(\left\|z\right\|^{3}\right)$. Invoking (2.7) and noting that $f\left(0,x,t\right)=0$ and $f_{z}\left(0,x,t\right)=0$, we can rewrite the previous equation, after some straightforward computations, as
$f\left(z,x,t\right)=\frac{1}{2}\mathbf{1}^{\top}\left(u_{xx}\left(x,t\right)\circ I_{x}zz^{\top}I_{x}\right)\mathbf{1}+\varepsilon\left(z,x,t\right)\qquad\text{as }\left\|z\right\|\rightarrow 0^{+},$ (3.1)
where $u_{xx}$ is the Hessian of $u$ with respect to $x$ and $I_{x}=\text{diag}\left(x^{\left(1\right)},x^{\left(2\right)}\right)$.
Next, let $R_{z}^{\mathbf{I}}$, $R_{z}^{\mathbf{II}}$ and $R_{z}^{\mathbf{III}}$ be three sets defined by
$R_{z}^{\mathbf{I}}$ $=\left\{z\in\mathbb{R}^{2}:\left\|z\right\|_{\infty}\leq z_{\max}^{\mathbf{I}}\right\},$
$R_{z}^{\mathbf{II}}$ $=\left\{z\in\mathbb{R}^{2}:z_{\max}^{\mathbf{I}}<\left\|z\right\|_{\infty}\leq z_{\max}^{\mathbf{II}}\right\},$
$R_{z}^{\mathbf{III}}$ $=\left\{z\in\mathbb{R}^{2}:z_{\max}^{\mathbf{II}}<\left\|z\right\|_{\infty}\leq z_{\max}^{\mathbf{III}}\right\},$
where $\left\|z\right\|_{\infty}=\max_{j=1,2}\left|z^{\left(j\right)}\right|$ and $0<z_{\max}^{\mathbf{I}}<z_{\max}^{\mathbf{II}}<z_{\max}^{\mathbf{III}}$ are given numbers, which will be specified in Section 4. The above three sets represent a partition of $R_{z}=\left\{z\in\mathbb{R}^{2}:\left\|z\right\|_{\infty}\leq z_{\max}^{\mathbf{III}}\right\}$, which is a square centered at the origin, as shown in Figure 1.
Figure 1: Partition of the integration domain $R_{z}$
For any given $N_{z}\in\mathbb{N}$, define a set of points $\mathbf{z}$ whose elements are
$z_{l_{1}l_{2}}=\left(\left(l_{1}+\frac{1}{2}\right)h_{z},\left(l_{2}+\frac{1}{2}\right)h_{z}\right)\qquad\left(l_{1},l_{2}=-N_{z},-N_{z}+1,\ldots,N_{z}-2,N_{z}-1\right),$
where $h_{z}=z_{\max}^{\mathbf{III}}/N_{z}$ denotes the mesh-width. Note that the point $z_{l_{1}l_{2}}$ is the center of the cell
$R_{l_{1}l_{2}}=\left[l_{1}h_{z},\left(l_{1}+1\right)h_{z}\right]\times\left[l_{2}h_{z},\left(l_{2}+1\right)h_{z}\right].$
We then consider the approximation
$\int_{\mathbb{R}_{*}^{2}}f\left(z,x,t\right)\ell\left(dz\right)\simeq\int_{R_{z}^{\mathbf{I}}}f\left(z,x,t\right)\ell\left(dz\right)+\int_{R_{z}^{\mathbf{II}}}f\left(z,x,t\right)\ell\left(dz\right)+\int_{R_{z}^{\mathbf{III}}}f\left(z,x,t\right)\ell\left(dz\right),$ (3.2)
where the individual terms on the right-hand side will be approximated in different ways: the first one will be transformed into a diffusion term by replacing the integrand function with its Taylor expansion; for the second one, a particular quadrature formula is used that takes into account the limiting singular behaviour of the Lévy measure as $\left\|z\right\|\rightarrow 0^{+}$; for the third one, a generic method is used.
By substituting (3.1) in the first integral in (3.2), it follows that
$\int_{R_{z}^{\mathbf{I}}}f\left(z,x,t\right)\ell\left(dz\right)\simeq\frac{1}{2}\mathbf{1}^{\top}\left(u_{xx}\left(x,t\right)\circ I_{x}\left(\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right)\right)I_{x}\right)\mathbf{1}.$ (3.3)
Here, the entries of the matrix $\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right)$ can be accurately approximated using a common numerical integrator.
Moving on to the second and third terms in (3.2), we consider a quadrature formula of the form
$\int_{R_{z}^{\mathbf{II}}\cup R_{z}^{\mathbf{III}}}f\left(z,x,t\right)\ell\left(dz\right)\simeq\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}f\left(z_{l_{1}l_{2}},x,t\right).$ (3.4)
Defining the coefficients
$\omega_{l_{1}l_{2}}=\begin{cases}0&\text{if }l_{1},l_{2}:z_{l_{1}l_{2}}\in R_{z}^{\mathbf{I}},\\ \left\|z_{l_{1}l_{2}}\right\|^{-2}\int_{R_{l_{1}l_{2}}}\left\|z\right\|^{2}\ell\left(dz\right)&\text{if }l_{1},l_{2}:z_{l_{1}l_{2}}\in R_{z}^{\mathbf{II}},\\ \ell\left(z_{l_{1}l_{2}}\right)h_{z}^{2}&\text{if }l_{1},l_{2}:z_{l_{1}l_{2}}\in R_{z}^{\mathbf{III}},\end{cases}$ (3.5)
a high level of accuracy is achieved, despite the integrand being singular at the origin. Clearly, the quadrature weights used in $R_{z}^{\mathbf{II}}$ are constructed as integrals of the Lévy measure, which turns out to be beneficial for the convergence behaviour (as $N_{z}\rightarrow\infty$). Analogously to the entries of the matrix $\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right)$ in (3.3), the integrals $\int_{R_{l_{1}l_{2}}}\left\|z\right\|^{2}\ell\left(dz\right)$ can be precomputed using a common numerical integrator. Regarding $R_{z}^{\mathbf{III}}$, the coefficients are obtained by applying the classical midpoint rule, see for example Quarteroni et al. (2007). Finally, note that the weights $\omega_{l_{1}l_{2}}$ are null over $R_{z}^{\mathbf{I}}$, as the first integral in (3.2) has already been approximated through (3.3).
Using (3.3) and substituting (2.7) into (3.4), we can define an approximating operator $\mathcal{A}_{\omega}$ and a number $r_{\omega}$ such that
$\mathcal{A}u\left(x,t\right)-ru\left(x,t\right)\simeq\mathcal{A}_{\omega}u\left(x,t\right)-r_{\omega}u\left(x,t\right)\qquad\text{for any }\left(x,t\right)\in\mathbb{R}_{\geq 0}^{2}\times\left(0,T\right],$
with
$\mathcal{A}_{\omega}u\left(x,t\right)$ $=\mu_{\omega}\left(x\right)^{\top}u_{x}\left(x,t\right)+\frac{1}{2}\mathbf{1}^{\top}\left(u_{xx}\left(x,t\right)\circ\Sigma_{\omega}\Sigma_{\omega}^{\top}\left(x\right)\right)\mathbf{1}+\left(\mathcal{B}_{\omega}u\right)\left(x,t\right)$ (3.6)
$r_{\omega}$ $=r+\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}$
where, for $i,j=1,2$,
$\mu_{\omega}^{\left(i\right)}\left(x\right)$ $=x^{\left(i\right)}\kappa_{\omega}^{\left(i\right)}$
$\kappa_{\omega}^{\left(i\right)}$ $=r-\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}\left(e^{z_{l_{1}l_{2}}^{\left(i\right)}}-1\right)$
$\left(\Sigma_{\omega}\Sigma_{\omega}^{\top}\right)^{\left(i,j\right)}\left(x\right)$ $=x^{\left(i\right)}x^{\left(j\right)}\left(\sigma_{\omega}\sigma_{\omega}^{\top}\right)^{\left(i,j\right)}$
$\sigma_{\omega}\sigma_{\omega}^{\top}$ $=\sigma\sigma^{\top}+\int_{R_{z}^{\mathbf{I}}}zz^{\top}\ell\left(dz\right)$
$\left(\mathcal{B}_{\omega}u\right)\left(x,t\right)$ $=\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}u\left(x+\gamma\left(z_{l_{1}l_{2}},x\right),t\right).$ (3.7)
Then, we approximate the solution $u$ of (2.8) by the function $v:\mathbb{R}_{\geq 0}^{2}\times\left[0,T\right]\rightarrow\mathbb{R}$ which solves the following problem
$\begin{cases}v_{t}\left(x,t\right)=\mathcal{A}_{\omega}v\left(x,t\right)-r_{\omega}v\left(x,t\right)&\text{for any }\left(x,t\right)\in\mathbb{R}_{\geq 0}^{2}\times\left(0,T\right]\\ v\left(x,0\right)=\phi\left(x\right).\end{cases}$ (3.8)
3.2 Spatial discretization
In this section, we successively consider the spatial discretization of the diffusion and summation terms in the operator $\mathcal{A}_{\omega}$. The convection term will be discussed in Section 3.3.
Let $R_{x}=\left[0,x_{\max}\right]\times\left[0,x_{\max}\right]$ be the truncated $x$-domain over which the solution to (3.8) is approximated and $N_{x}\in\mathbb{N}$ be a given number of spatial grid points. Here, $x_{\max}$ is heuristically chosen in such a way that the localization error is negligible. We construct a spatial grid $\mathbf{x}$ in $R_{x}$ by applying, in each dimension, a strictly increasing and smooth transformation $\varphi$ to an artificial uniform grid. Let
$x_{m}=\varphi\left(\varphi^{-1}\left(0\right)+\frac{\varphi^{-1}\left(x_{\max}\right)-\varphi^{-1}\left(0\right)}{N_{x}}m\right)\qquad\left(m=0,1,\ldots,N_{x}\right)$
with $h_{x,m}=x_{m}-x_{m-1}$. The elements of $\mathbf{x}$ are defined by
$x_{m_{1}m_{2}}=\left(x_{m_{1}},x_{m_{2}}\right)\qquad\left(m_{1},m_{2}=0,1,\ldots,N_{x}\right).$
The function $\varphi$ will be chosen in such a way that relatively many points are placed in a region of financial and numerical interest.
In what follows, we denote the values over $\mathbf{x}$ of any given function $g:R_{x}\times\left[0,T\right]\rightarrow\mathbb{R}$ by the vector
$g\left(\mathbf{x},t\right)=\left[g\left(x_{00},t\right),g\left(x_{10},t\right),\ldots,g\left(x_{N_{x}-1,N_{x}},t\right),g\left(x_{N_{x}N_{x}},t\right)\right]^{\top}.$ (3.9)
3.2.1 Diffusion term
In this subsection, we construct a semi-discrete diffusion matrix $D$ such that
$Dv\left(\mathbf{x},t\right)\simeq\left[\frac{1}{2}\mathbf{1}^{\top}\left(v_{xx}\left(x_{m_{1}m_{2}},t\right)\circ\Sigma_{\omega}\Sigma_{\omega}^{\top}\left(x_{m_{1}m_{2}}\right)\right)\mathbf{1}\right]_{m_{1},m_{2}=0,1,\ldots,N_{x}},$ (3.10)
where the right-hand side is a vector, whose elements are ordered according to (3.9).
To this purpose, in each spatial dimension, we approximate the first- and second-order derivatives of a given smooth function $g:\mathbb{R}\rightarrow\mathbb{R}$ by the following second-order central finite difference schemes
$g^{\prime}\left(x_{m}\right)$ $\simeq\alpha_{m}^{\left(-1\right)}g\left(x_{m-1}\right)+\alpha_{m}^{\left(0\right)}g\left(x_{m}\right)+\alpha_{m}^{\left(1\right)}g\left(x_{m+1}\right)$
$g^{\prime\prime}\left(x_{m}\right)$ $\simeq\beta_{m}^{\left(-1\right)}g\left(x_{m-1}\right)+\beta_{m}^{\left(0\right)}g\left(x_{m}\right)+\beta_{m}^{\left(1\right)}g\left(x_{m+1}\right)$
with coefficients
$\alpha_{m}^{\left(-1\right)}=\frac{-h_{x,m+1}}{h_{x,m}\left(h_{x,m}+h_{x,m+1}\right)},\qquad\alpha_{m}^{\left(0\right)}=\frac{h_{x,m+1}-h_{x,m}}{h_{x,m}h_{x,m+1}},\qquad\alpha_{m}^{\left(1\right)}=\frac{h_{x,m}}{h_{x,m+1}\left(h_{x,m}+h_{x,m+1}\right)},$
$\beta_{m}^{\left(-1\right)}=\frac{2}{h_{x,m}\left(h_{x,m}+h_{x,m+1}\right)},\qquad\beta_{m}^{\left(0\right)}=\frac{-2}{h_{x,m}h_{x,m+1}},\qquad\beta_{m}^{\left(1\right)}=\frac{2}{h_{x,m+1}\left(h_{x,m}+h_{x,m+1}\right)}.$
Concerning the boundary of the truncated spatial domain, we modify the previous formulae in the following way. At the lower boundary $x_{0}=0$, the first- and second-order derivative terms in (3.8) vanish. Hence, it is natural to choose $\alpha_{0}^{\left(j\right)}=0$ and $\beta_{0}^{\left(j\right)}=0$ for any $j=\left\{-1,0,1\right\}$. At the upper boundary $x_{N_{x}}=x_{\max}$, we make the natural assumption that the solution $v$ behaves linearly in $x$, thus we choose $\beta_{N_{x}}^{\left(j\right)}=0$ for any $j=\left\{-1,0,1\right\}$, and we approximate the first-order derivative by the first-order backward finite difference scheme.
Noting that $\mathbf{x}$ is the Cartesian product of two identical 1-dimensional grids, by employing the 1-directional finite difference formulae in both the spatial dimensions, it leads to the matrix $D$ defined by
$D=\frac{1}{2}\left(\sigma_{\omega}\sigma_{\omega}^{\top}\right)^{\left(1,1\right)}I\otimes I_{\mathbf{x}}^{2}D_{2}+\left(\sigma_{\omega}\sigma_{\omega}^{\top}\right)^{\left(1,2\right)}I_{\mathbf{x}}D_{1}\otimes I_{\mathbf{x}}D_{1}+\frac{1}{2}\left(\sigma_{\omega}\sigma_{\omega}^{\top}\right)^{\left(2,2\right)}I_{\mathbf{x}}^{2}D_{2}\otimes I.$ (3.11)
Here, $I\in\mathbb{R}^{\left(N_{x}+1\right)\times\left(N_{x}+1\right)}$ is the identity matrix, $I_{\mathbf{x}}=\text{diag}\left(x_{0}^{\left(i\right)},\ldots,x_{N_{x}}^{\left(i\right)}\right)$ and $\otimes$ denotes the Kronecker product. [^4] In this paper, we use the convention $AB\otimes CD=\left(AB\right)\otimes\left(CD\right)$, for any suitable matrices $A,B,C,D$. The matrices $D_{1},D_{2}\in\mathbb{R}^{\left(N_{x}+1\right)\times\left(N_{x}+1\right)}$ are the matrices representing numerical differentiation of first- and second-order by the relevant finite difference formulae above. The mixed derivative has been approximated by applying the finite difference formula for the first-order derivative subsequently in the two spatial dimensions.
3.2.2 Summation term
In this section, we derive an efficient method to approximate the summation term $\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{x},t\right)$ given the values of $v\left(\mathbf{x},t\right)$. Unlike the differential component of $\mathcal{A}_{\omega}$, we do not construct a matrix $B_{\omega}$ such that $\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{x},t\right)\simeq B_{\omega}v\left(\mathbf{x},t\right)$, as this matrix would be large and dense.
Assuming that the values of $v$ are known for all $\left(x,t\right)\in R_{x}\times\left[0,T\right]$, using formula (3.7) to directly evaluate $\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{x},t\right)$ would require $O\left(N_{x}^{2}N_{z}^{2}\right)$ elementary operations, which is computationally too expensive. For this reason, a particularly efficient method combining interpolation and FFT is considered, which extends the approach by Wang et al. (2007).
Let $N_{y}^{-},N_{y}^{+}\in\mathbb{N}$ be any given natural numbers and let $\mathbf{y}^{\text{out}}$ and $\mathbf{y}^{\text{in}}$ be two grids of points defined by [^5] The superscripts stand for “input” and “output”.
$y_{m_{1}m_{2}}^{\text{out}}$ $=\left(e^{m_{1}h_{z}},e^{m_{2}h_{z}}\right)\qquad\left(m_{1},m_{2}=-N_{y}^{-},-N_{y}^{-}+1,\ldots,N_{y}^{+}-1,N_{y}^{+}\right),$
$y_{m_{1}m_{2}}^{\text{in}}$ $=\left(e^{\left(m_{1}+\frac{1}{2}\right)h_{z}},e^{\left(m_{2}+\frac{1}{2}\right)h_{z}}\right)\qquad\left(m_{1},m_{2}=-N_{z}-N_{y}^{-},-N_{z}-N_{y}^{-}+1,\ldots,N_{z}+N_{y}^{+}-2,N_{z}+N_{y}^{+}-1\right),$
then it holds that
$\left(\mathcal{B}_{\omega}v\right)\left(y_{m_{1}m_{2}}^{\text{out}},t\right)=\sum_{l_{1},l_{2}=-N_{z}}^{N_{z}-1}\omega_{l_{1}l_{2}}v\left(y_{l_{1}+m_{1},l_{2}+m_{2}}^{\text{in}},t\right)\qquad\left(m_{1},m_{2}=-N_{y}^{-},-N_{y}^{-}+1,\ldots,N_{y}^{+}-1,N_{y}^{+}\right).$ (3.12)
Clearly, the summation term (3.12) can be viewed as a discrete 2-dimensional cross-correlation. It is well known, see for instance Plonka et al. (2018, Chapter 3), that it can be written in the form
$\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{y}^{\text{out}},t\right)=\tilde{I}Cv\left(\mathbf{y}^{\text{in}},t\right)$ (3.13)
where:
  • $\bullet$
    $C\in\mathbb{R}^{\left(\sharp\text{in}\right)^{2}\times\left(\sharp\text{in}\right)^{2}}$ is a circulant matrix whose first row is given by $C_{1,\cdot}^{\top}$ with
    $C_{1,\cdot}=\text{vec}\left(\left[\begin{array}[]{cc}\Omega&0_{\sharp\mathbf{z}\times\left(\sharp\text{in}-\sharp\mathbf{z}\right)}\\ 0_{\left(\sharp\text{in}-\sharp\mathbf{z}\right)\times\sharp\mathbf{z}}&0_{\left(\sharp\text{in}-\sharp\mathbf{z}\right)\times\left(\sharp\text{in}-\sharp\mathbf{z}\right)}\end{array}\right]\right).$ (3.14)
    Here, $0_{P\times Q}$ denotes the null matrix of dimensions $P\times Q$, $\text{vec}\left(\cdot\right)$ denotes the vectorization of a matrix, $\sharp$ indicates the number of points of a given grid in one direction and $\Omega\in\mathbb{R}^{\sharp\mathbf{z}\times\sharp\mathbf{z}}$ is the matrix whose entries are the coefficients $\omega_{l_{1}l_{2}}$ defined by (3.5). For an example of a matrix $C$, we refer to Appendix B.
    The matrix-vector multiplication $Ca$, for any given vector $a\in\mathbb{R}^{\left(\sharp\text{in}\right)^{2}\times 1}$, can be obtained by two (1-dimensional) FFTs and one (1-dimensional) inverse FFT, requiring just $O\left(\left(\sharp\text{in}\right)^{2}\cdot\log\sharp\text{in}\right)$ elementary operations. The pertinent formula is
    $Ca=\text{ifft}\left(\text{fft}\left(C_{1,\cdot}\right)^{H}\circ\text{fft}\left(a\right)\right),$ (3.15)
    where H denotes the complex conjugate.
  • $\bullet$
    $\tilde{I}\in\mathbb{R}^{\left(\sharp\text{out}\right)^{2}\times\left(\sharp\text{in}\right)^{2}}$ is obtained from the identity matrix $I\in\mathbb{R}^{\left(\sharp\text{in}\right)^{2}\times\left(\sharp\text{in}\right)^{2}}$ by removing the rows corresponding to the zeros in the following vector
    $\text{vec}\left(\left[\begin{array}[]{cc}1_{\sharp\text{out}\times\sharp\text{out}}&0_{\sharp\text{out}\times\left(\sharp\text{in}-\sharp\text{out}\right)}\\ 0_{\left(\sharp\text{in}-\sharp\text{out}\right)\times\sharp\text{out}}&0_{\left(\sharp\text{in}-\sharp\text{out}\right)\times\left(\sharp\text{in}-\sharp\text{out}\right)}\end{array}\right]\right).$
    Here, $1_{P\times P}$ denotes a $P\times P$ matrix whose elements are all equal to 1. We note that the matrix-vector multiplication $Cv\left(\mathbf{y}^{\text{in}},t\right)$ in (3.13) returns a value also for grid points that can be discarded. The purpose of $\tilde{I}$ is precisely to extract only those entries that correspond to $\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{y}^{\text{out}},t\right)$.
In order to obtain an approximation to $\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{x},t\right)$ using (3.13), we need to interpolate both the input and the output value in (3.13) since $\mathbf{y}^{\text{in}}$ and $\mathbf{y}^{\text{out}}$ are generally different from $\mathbf{x}$. Let $T^{\text{in}}\in\mathbb{R}^{\left(\sharp\text{in}\right)^{2}\times\left(N_{x}+1\right)^{2}}$ be a matrix representing an interpolation procedure from the $\mathbf{x}$ grid to the $\mathbf{y}^{\text{in}}$ grid and let $T^{\text{out}}\in\mathbb{R}^{\left(N_{x}+1\right)^{2}\times\left(\sharp\text{in}\right)^{2}}$ be a matrix representing an interpolation procedure from the $\mathbf{y}^{\text{out}}$ grid to the $\mathbf{x}$ grid. Then
$v\left(\mathbf{y}^{\text{in}},t\right)$ $\simeq T^{\text{in}}v\left(\mathbf{x},t\right),$ (3.16)
$\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{x},t\right)$ $\simeq T^{\text{out}}\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{y}^{\text{out}},t\right).$ (3.17)
Note that, by using Lagrange interpolation, the interpolation matrices are sparse and have at most $P+1$ nonzero entries per row, where $P$ is the polynomial degree. Let $M$ be the number of rows, it follows that the corresponding matrix–vector multiplications require a number of operations of order $O\left(MP\right)$, and are therefore negligible compared with multiplication performed via FFT.
From (3.13), (3.16) and (3.17), we arrive at the approximation
$\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{x},t\right)\simeq B_{\omega}v\left(\mathbf{x},t\right),$ (3.18)
where $B_{\omega}\in\mathbb{R}^{\left(N_{x}+1\right)^{2}\times\left(N_{x}+1\right)^{2}}$ is given by
$B_{\omega}=T^{\text{out}}\tilde{I}CT^{\text{in}}.$ (3.19)
We emphasize that $B_{\omega}$ is only used for notational purposes and never explicitly computed. To compute the right-hand side of (3.18), we always use
$B_{\omega}V=T^{\text{out}}\tilde{I}\,\text{ifft}\left(\text{fft}\left(C_{1,\cdot}\right)^{H}\circ\text{fft}\left(T^{\text{in}}V\right)\right),$ (3.20)
for any vector $V\in\mathbb{R}^{\left(N_{x}+1\right)^{2}\times 1}$. Figure
Figure 2: Diagram of the scheme used to approximate $\left(\mathcal{B}_{\omega}v\right)\left(\mathbf{x},t\right)$
[…]
$\begin{array}[]{c}\text{interpolation}\\ \text{from $\mathbf{x}$ to }\mathbf{y}^{\text{in}}\end{array}$ $\downarrow$ $\uparrow$ $\begin{array}[]{c}\text{interpolation}\\ \text{from $\mathbf{y}^{\text{out}}$ to }\mathbf{x}\end{array}$
[…]
$\begin{array}[]{c}\text{matrix-vector}\\ \text{multiplication by FFT}\end{array}$
2 provides a schematic illustration of how FFT and interpolation are combined to evaluate (3.20).
3.2.3 Cell averaging
We conclude the spatial discretization with a technique for handling the non-smoothness of the initial function $\phi$ of (2.8). As it turns out, pointwise valuation of $\phi$ over the spatial grid can lead to deteriorated (spatial) convergence behaviour, which can be alleviated by applying cell averaging.
Let
$x_{m+\frac{1}{2}}$ $=\frac{1}{2}\left(x_{m}+x_{m+1}\right)\qquad\left(m=0,1,\ldots,N_{x}-1\right)$
$h_{x,m+\frac{1}{2}}$ $=x_{m+\frac{1}{2}}-x_{m-\frac{1}{2}}\qquad\left(m=0,1,\ldots,N_{x}\right)$
with $x_{-\frac{1}{2}}=-x_{\frac{1}{2}}$ and $x_{N_{x}+\frac{1}{2}}=2x_{\max}-x_{N_{x}-\frac{1}{2}}$. Then, we use the approximation
$v\left(x_{m_{1}m_{2}},0\right)\simeq\frac{1}{h_{x,m_{1}+\frac{1}{2}}h_{x,m_{2}+\frac{1}{2}}}\int_{x_{m_{1}-\frac{1}{2}}}^{x_{m_{1}+\frac{1}{2}}}\int_{x_{m_{2}-\frac{1}{2}}}^{x_{m_{2}+\frac{1}{2}}}\phi\left(x_{1},x_{2}\right)dx_{2}dx_{1},$ (3.21)
whenever the cell $\left[x_{m_{1}-\frac{1}{2}},x_{m_{1}+\frac{1}{2}}\right)\times\left[x_{m_{2}-\frac{1}{2}},x_{m_{2}+\frac{1}{2}}\right)$ has a nonempty intersection with the set of non-smoothness of $\phi$.
3.3 Temporal discretization: the semi-Lagrangian $\theta$-method
The problem (3.8) can be convection-dominated. To account for this, we shall consider temporal discretization using the $\theta$-method combined with the semi-Lagrangian approach, as described by Spiegelman & Katz (2006). The semi-Lagrangian method follows, in each time step, the characteristics backwards in time to determine the departure points of the spatial grid points.
Let $x:\left[0,T\right]\rightarrow\mathbb{R}_{\geq 0}^{2}$ and $v^{*}:\left[0,T\right]\rightarrow\mathbb{R}$ such that $v^{*}\left(t\right)=v\left(x\left(t\right),t\right)$. The derivative of $v^{*}$ is given by
$v_{t}^{*}\left(t\right)=v_{t}\left(x\left(t\right),t\right)+x_{t}\left(t\right)^{\top}v_{x}\left(x\left(t\right),t\right).$
Assume $x$ satisfies the following (linear) ODE:
$x_{t}\left(t\right)=-\mu_{\omega}\left(x\left(t\right)\right)\qquad\left(0<t\leq T\right).$ (3.22)
Then
$v_{t}^{*}\left(t\right)=\left(\mathcal{A}_{\omega}^{\text{SL}}-r_{\omega}\right)v\left(x\left(t\right),t\right)\qquad\left(0<t\leq T\right),$ (3.23)
where
$\mathcal{A}_{\omega}^{\text{SL}}v\left(x,t\right)=\frac{1}{2}\mathbf{1}^{\top}\left(v_{xx}\left(x,t\right)\circ\Sigma_{\omega}\Sigma_{\omega}^{\top}\left(x\right)\right)\mathbf{1}+\left(\mathcal{B}_{\omega}v\right)\left(x,t\right).$
Clearly, $\mathcal{A}_{\omega}^{\text{SL}}$ is obtained from $\mathcal{A}_{\omega}$ by omitting the convection term.
Let parameter $\theta\in\left[0,1\right]$. Let $\mathbf{t}=\left(t_{n}\right)_{n=0}^{N_{t}}$ be any given uniform grid with step size $h_{t}=\frac{T}{N_{t}}$. For any given $n=1,2,\ldots,N_{t}$, approximating (3.23) using the $\theta$-method and substituting the definition of $v^{*}$, we obtain
$\frac{v\left(x\left(t_{n}\right),t_{n}\right)-v\left(x\left(t_{n-1}\right),t_{n-1}\right)}{h_{t}}\simeq\theta\left(\mathcal{A}_{\omega}^{\text{SL}}-r_{\omega}\right)v\left(x\left(t_{n}\right),t_{n}\right)+\left(1-\theta\right)\left(\mathcal{A}_{\omega}^{\text{SL}}-r_{\omega}\right)v\left(x\left(t_{n-1}\right),t_{n-1}\right).$ (3.24)
The approximation (3.24) holds along any trajectory satisfying (3.22). In each given time step from $t_{n-1}$ to $t_{n}$, the semi-Lagrangian approach involves selecting the set of trajectories that intersect the points $\left(\mathbf{x},t_{n}\right)$, ensuring that an approximation is defined on the fixed grid $\mathbf{x}$. Let $\mathbf{x}^{\text{SL}}$ denote the grid corresponding to $t_{n-1}$ along this set of trajectories. Its elements are given by $x_{m_{1}m_{2}}^{\text{SL}}=\left(x_{m_{1}}^{\text{SL}},x_{m_{2}}^{\text{SL}}\right)$ where $x_{m_{i}}^{\text{SL}}$ is obtained by (3.22) as
$x_{m_{i}}^{\text{SL}}=x_{m_{i}}e^{\kappa_{\omega}^{\left(i\right)}h_{t}}\qquad\left(m_{i}=0,1,\ldots,N_{x}\right).$
Then (3.24) becomes
$\frac{v\left(\mathbf{x},t_{n}\right)-v\left(\mathbf{x}^{\text{SL}},t_{n-1}\right)}{h_{t}}\simeq\theta\left(\mathcal{A}_{\omega}^{\text{SL}}-r_{\omega}\right)v\left(\mathbf{x},t_{n}\right)+\left(1-\theta\right)\left(\mathcal{A}_{\omega}^{\text{SL}}-r_{\omega}\right)v\left(\mathbf{x}^{\text{SL}},t_{n-1}\right).$
Interpolation is employed to acquire approximations at the grid $\mathbf{x}^{\text{SL}}$. Let $T^{\text{SL}}\in\mathbb{R}^{\left(N_{x}+1\right)^{2}\times\left(N_{x}+1\right)^{2}}$ be a matrix representing an interpolation procedure from the $\mathbf{x}$ grid to the $\mathbf{x}^{\text{SL}}$ grid. Together with the approximation of the diffusion and summation terms, discussed in Section 3.2, we obtain
$v\left(\mathbf{x}^{\text{SL}},t_{n-1}\right)$ $\simeq T^{\text{SL}}v\left(\mathbf{x},t_{n-1}\right),$ (3.25)
$\left(\mathcal{A}_{\omega}^{\text{SL}}-r_{\omega}\right)v\left(\mathbf{x},t_{n}\right)$ $\simeq\left(D+B_{\omega}-r_{\omega}I\right)v\left(\mathbf{x},t_{n}\right),$
$\left(\mathcal{A}_{\omega}^{\text{SL}}-r_{\omega}\right)v\left(\mathbf{x}^{\text{SL}},t_{n-1}\right)$ $\simeq T^{\text{SL}}\left(D+B_{\omega}-r_{\omega}I\right)v\left(\mathbf{x},t_{n-1}\right).$
This leads to the following natural definition of the approximation $V^{n}$ to the exact solution vector $v\left(\mathbf{x},t_{n}\right)$:
$\left(I-h_{t}\theta\left(D+B_{\omega}-r_{\omega}I\right)\right)V^{n}=T^{\text{SL}}\left(I+h_{t}\left(1-\theta\right)\left(D+B_{\omega}-r_{\omega}I\right)\right)V^{n-1}$ (3.26)
for $n=1,2,\ldots,N_{t}$. The initial vector $V^{0}$ is defined by pointwise valuation on the spatial grid $\mathbf{x}$ of the pay-off function $\phi$, except near the set of non-smoothness, where cell averaging is employed (see Section 3.2.3). The time-stepping scheme (3.26) is called the semi-Lagrangian $\theta$-method. We shall apply (3.26) with $\theta=\frac{1}{2}$, which is also called the semi-Lagrangian Crank–Nicolson method. Here, to account for the non-smoothness of $\phi$, a damping procedure is used where the first time step (i.e. $n=1$) is replaced by four time steps of size equal to $\frac{1}{4}h_{t}$ of (3.26) with $\theta=1$.
It remains to consider the treatment of the discretized integral term in (3.26), represented formally by the matrix $B_{\omega}$. Recall from Section 3.2.2 that $B_{\omega}$ is never actually computed. To effectively handle this term, we shall employ fixed-point iteration:
$\left(I-h_{t}\theta\left(D-r_{\omega}I\right)\right)Y^{n,k}=h_{t}\theta B_{\omega}Y^{n,k-1}+T^{\text{SL}}\left(I+h_{t}\left(1-\theta\right)\left(D-r_{\omega}I\right)\right)V^{n-1}+h_{t}\left(1-\theta\right)T^{\text{SL}}B_{\omega}V^{n-1}$ (3.27)
for $k=1,2,\ldots$. Here matrix-vector multiplications involving $B_{\omega}$ are always computed by the efficient FFT algorithm of Section 3.2.2. For a given tolerance $tol>0$ sufficiently small, we use the following stopping criterion
$\max_{m_{1},m_{2}}\frac{\left|Y_{m_{1}m_{2}}^{n,k}-Y_{m_{1}m_{2}}^{n,k-1}\right|}{\max\left\{1,\left|Y_{m_{1}m_{2}}^{n,k}\right|\right\}}<tol$ (3.28)
and define $V^{n}=Y^{n,k}$.
The starting vector $Y^{n,0}$ for the fixed-point iteration is commonly chosen in the literature as $Y^{n,0}=V^{n-1}$. Here, we shall consider a more accurate starting vector, defined by higher-order extrapolation from known approximations at previous temporal grid points:
$Y^{n,0}=\begin{cases}V^{n-1}&n=1,\\ 2V^{n-1}-V^{n-2}&n=2,\\ 3V^{n-1}-3V^{n-2}+V^{n-3}&n=3,\\ 4V^{n-1}-6V^{n-2}+4V^{n-3}-V^{n-4}&n\geq 4.\end{cases}$ (3.29)
This yields a significant reduction in the number of fixed-point iterations compared to the common choice.
Finally, for the linear system in (3.27) we apply the BiCGSTAB iterative solver using an ILU preconditioner.
Our complete algorithm for the numerical solution of problem (2.8) is outlined in Algorithm 1.
Algorithm 1 Outline of the algorithm
precomputations:
$\bullet$
define the grids $\mathbf{z}$, $\mathbf{x}$, $\mathbf{y}^{\text{in}}$, $\mathbf{y}^{\text{out}}$, $\mathbf{t}$ and $\mathbf{x}^{\text{SL}}$
$\bullet$
define the matrix $D$ given by (3.11) and compute the ILU factorization of $I-h_{t}\theta\left(D-r_{\omega}I\right)$
$\bullet$
define the vector $C_{1,\cdot}$ given by (3.14) and compute $\text{fft}\left(C_{1,\cdot}\right)$
$\bullet$
define the matrices $T^{\text{in}}$, $T^{\text{out}}$ and $T^{\text{SL}}$ given by (3.16), (3.17) and (3.25)
$\bullet$
choose $\theta=\frac{1}{2}$
time-stepping:
compute $V^{0}=\phi\left(\mathbf{x}\right)$ and apply cell averaging (3.21)
for $n=1,2,\ldots,N_{t}$
1.
compute $B_{\omega}V^{n-1}$ using (3.20)
2.
compute $W^{n-1}=T^{\text{SL}}\left(I+h_{t}\left(1-\theta\right)\left(D-r_{\omega}I\right)\right)V^{n-1}+h_{t}\left(1-\theta\right)T^{\text{SL}}B_{\omega}V^{n-1}$
3.
compute $Y^{n,0}$ given by (3.29)
4.
for $k=1,2,\ldots$
i.
compute $B_{\omega}Y^{n,k-1}$ using (3.20)
ii.
solve $\left(I-h_{t}\theta\left(D-r_{\omega}I\right)\right)Y^{n,k}=h_{t}\theta B_{\omega}Y^{n,k-1}+W^{n-1}$ using BiCGSTAB
5.
end for if $Y^{n,k}$ satisfies (3.28)
6.
let $V^{n}=Y^{n,k}$
end for
4 Numerical experiments
We consider an European put-on-the-average option, which has the pay-off function
$\phi\left(x\right)=\max\left(K-\frac{1}{2}\left(x^{\left(1\right)}+x^{\left(2\right)}\right),0\right)$
with fixed strike price $K>0$. Clearly, $\phi$ is non-smooth over the set $\left\{x\in\mathbb{R}_{\geq 0}^{2}:x^{\left(1\right)}+x^{\left(2\right)}=2K\right\}$. To define the non-uniform grid $\mathbf{x}$, we use the same transformation $\varphi$ as in in ’t Hout & Lamotte (2023). Let $c$, $x_{\rm int}$ be two given positive numbers. We choose the function $\varphi$ in Section 3.2 as
$\varphi\left(\xi\right)=\begin{cases}c\xi&0\leq\xi\leq\xi_{\rm int},\\ x_{\rm int}+c\sinh\left(\xi-\xi_{\rm int}\right)&\xi_{\rm int}<\xi\leq\xi_{\rm max},\end{cases}$
with
$\xi_{\rm int}=\frac{x_{\rm int}}{c},\quad\xi_{\rm max}=\xi_{\rm int}+\sinh^{-1}\left(\frac{x_{\rm max}-x_{\rm int}}{c}\right).$
In this way, the resulting spatial grid in each direction is uniform over $\left[0,x_{\rm int}\right]$, whereas in $\left[x_{\rm int},x_{\rm max}\right]$ the distance between consecutive grid points gradually increases as one moves away from $x_{\rm int}$. The limit of the fraction of spatial grid points within the interval $\left[0,x_{\rm int}\right]$ as $N_{x}\rightarrow\infty$, denoted by $F$, is given by
$F=\frac{\xi_{\rm int}}{\xi_{\rm max}}=\left(1+\frac{c}{x_{\rm int}}\sinh^{-1}\left(\frac{x_{\rm max}-x_{\rm int}}{c}\right)\right)^{-1}.$
Note that $F\rightarrow\frac{x_{\rm int}}{x_{\rm max}}$ as $c\rightarrow\infty$, which corresponds to the uniform case.
Moving on to the Lévy measure, we model the jump component in (2.1) by a pure-jump 2-dimensional Normal Tempered Stable process. It is characterized by the parameters $0\leq\alpha<1$, $\delta>0$, $\lambda>0$, $\eta\in\mathbb{R}^{2\times 1}$ and a positive definite symmetric matrix $\rho\in\mathbb{R}^{2\times 2}$. The case where $\alpha=0$ is known as Variance Gamma, while the case where $\alpha=\frac{1}{2}$ is known as Normal Inverse Gaussian. Both are commonly used to model financial dynamics. The Lévy measure is given by
$\ell\left(z\right)=\frac{\delta}{\pi}\sqrt{\frac{\left(\left\|\eta\right\|_{\rho}^{2}+2\lambda\right)^{1+\alpha}}{\det\left[\rho\right]}}K_{1+\alpha}\left(\sqrt{\left\|\eta\right\|_{\rho}^{2}+2\lambda}\left\|z\right\|_{\rho}\right)\left\|z\right\|_{\rho}^{-1-\alpha}e^{\left\langle\eta,z\right\rangle_{\rho}}$
where $K_{\nu}\left(\tau\right)=\frac{1}{2}\int_{0}^{\infty}y^{\nu-1}e^{-\frac{1}{2}\tau\left(y+y^{-1}\right)}dy$, for $\tau>0$, denotes the modified Bessel function of the second kind, [^6] See Schoutens (2003, Appendix A). $\left\langle x,y\right\rangle_{\rho}=x^{\top}\rho^{-1}y$ and $\left\|x\right\|_{\rho}=\sqrt{\left\langle x,x\right\rangle_{\rho}}$ is its induced norm. The constants $A_{\ell},B_{\ell}$ and $C_{\ell}$ in (2.5) are defined, with respect to $\left\|\cdot\right\|_{\rho}$, as
$\left(A_{\ell},B_{\ell},C_{\ell}\left(h\right)\right)=\left(2\alpha,\sqrt{\left\|\eta\right\|_{\rho}^{2}+2\lambda}-\left\|\eta\right\|_{\rho},\frac{2^{\alpha}\delta\Gamma\left(1+\alpha\right)}{\pi\sqrt{\det\left[\rho\right]}}e^{h\left\|\eta\right\|_{\rho}}\right).$
The variance of the random variable $L\left(t\right)=\int_{0}^{t}\int_{\mathbb{R}_{*}^{2}}z\tilde{\Pi}\left(dt,dz\right)$, for $t\in\left[0,T\right]$, is given by
$\mathbb{V}\left[L\left(t\right)\right]=t\cdot\delta\frac{\Gamma\left(2-\alpha\right)}{\lambda^{2-\alpha}}\left(\rho\lambda^{1-\alpha}+\eta\eta^{\top}\right).$
We refer to Appendix A for further details.
Table 1
Table 1: Parameter sets
Parameters VG0 VG1 NIG0 NIG1
$\alpha$ 0 0 $\frac{1}{2}$ $\frac{1}{2}$
$\lambda$ 1 6 20766.4 57.1108
$\delta$ 1 6 0.77576 4.26367
$\eta^{\left(1\right)}$ -0.1 -0.1 -37.688 -0.295846
$\eta^{\left(2\right)}$ -0.2 -0.2 -2.224 -0.292984
$\rho^{\left(1,1\right)}$ 0.09 0.01 3.984 0.037021
$\rho^{\left(1,2\right)}$ 0.06 0 3.160 0.026574
$\rho^{\left(2,2\right)}$ 0.16 0.0225 3.512 0.054613
$r$ 0.05 0 0 0
$T$ 1 $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}$
$K$ 100 100 100 100
lists four sets of representative parameter values where we always take the diffusion matrix $\sigma$ equal to zero. Table 2
Table 2: Standard deviation and correlation coefficient
VG0 VG1 NIG0 NIG1
$\sqrt{\mathbb{V}\left[L^{\left(1\right)}\left(1\right)\right]}$ 0.3162 0.1080 0.1958 0.1943
$\sqrt{\mathbb{V}\left[L^{\left(2\right)}\left(1\right)\right]}$ 0.4472 0.1707 0.1830 0.2352
$\frac{\textrm{cov}\left[L^{\left(1\right)}\left(1\right),L^{\left(2\right)}\left(1\right)\right]}{\sqrt{\mathbb{V}\left[L^{\left(1\right)}\left(1\right)\right]\mathbb{V}\left[L^{\left(2\right)}\left(1\right)\right]}}$ 0.5656 0.1807 0.8417 0.5975
contains the corresponding standard deviations and correlation coefficients. The sets VG0 and NIG0 are taken from Hilber et al. (2013, page 208) and Rydberg (1997, Figure 8), respectively. The VG1 set was designed by us based on VG0. Finally, the NIG1 set was obtained via standard maximum likelihood estimation [^7] The density function for the case where $\alpha\in\left\{0,\frac{1}{2}\right\}$ can be found in Appendix A. using the close price data of S&P500 (^GSPC) and EUROSTOXX50 (^STOXX50E), retrieved from Yahoo Finance, covering the period from 01/01/2014 to 31/12/2024. In particular, we apply the methodology used by Hainaut & Le Courtois (2014) to the logarithmic return of the price, i.e. $d\ln X$.
The following list specifies all choices for the values of the parameters of our numerical scheme:
  • $\bullet$
    $N_{z}=2N_{x}$ and $N_{t}=\text{round}\left[\frac{1}{2}N_{x}\right]$. Clearly, with this choice, the three mesh widths are directly proportional to each other.
  • $\bullet$
    $z_{\max}^{\mathbf{I}}=2h_{z}$. This choice is motivated by the fact that the artificial diffusion acts over a small region around the origin.
  • $\bullet$
    $z_{\max}^{\mathbf{II}}=\sqrt{0.1}z_{\max}^{\mathbf{III}}$. In this way, the size of $R_{z}^{\mathbf{II}}$ is about 10% of the full integration domain $R_{z}$.
  • $\bullet$
    $z_{\max}^{\mathbf{III}}=\max\left\{\left\|z\right\|_{\infty}:z\in\mathbb{R}^{2},\ell\left(z\right)=10^{-8}\right\}$. Since the Lévy measure decays at least exponentially as $\left\|z\right\|\rightarrow\infty$, we ensure that $\ell\left(z\right)<10^{-8}$ for all $z\in\mathbb{R}^{2}$ such that $\left\|z\right\|_{\infty}>z_{\max}^{\mathbf{III}}$.
  • $\bullet$
    $x_{\rm int}=\frac{5}{2}K$. The non-smoothness set of $\phi$ is contained in the portion of $R_{x}$ where the grid $\mathbf{x}$ is uniform.
  • $\bullet$
    $x_{\rm max}$ was heuristically chosen as $57K$ for VG0, $5K$ for VG1, $6K$ for NIG0, and $7K$ for NIG1.
  • $\bullet$
    $c$ is chosen such that $F=\max\left(65\%,\frac{x_{\rm int}}{x_{\rm max}}\right)$. In this way, approximately at least 65% of the spatial grid points in each given direction are in the interval $\left[0,x_{\rm int}\right]$.
  • $\bullet$
    $N_{y}^{-}=\text{ceil}\left[-\frac{1}{h_{z}}\ln\left(x_{1}\right)\right]+N_{y}^{*}$ and $N_{y}^{+}=\text{ceil}\left[\frac{1}{h_{z}}\ln\left(x_{\rm max}\right)\right]+N_{y}^{*}$ for some given $N_{y}^{*}\in\mathbb{N}_{0}$. This choice minimises the need for extrapolation in (3.17) as it is necessary to extrapolate just to the grid points $x_{m_{1}m_{2}}$ with either $m_{1}=0$ or $m_{2}=0$. This is done in a linear fashion. In (3.16), we set $v\left(x,t\right)=0$ whenever $x\notin R_{x}$.
  • $\bullet$
    $N_{y}^{*}$ is taken as the minimal $n\in\mathbb{N}_{0}$ such that the maximal prime factor of $\sharp\textrm{in}=N_{y}^{-}+N_{y}^{+}+2N_{z}$ is at most 7. This is beneficial for the efficiency of the FFT.
  • $\bullet$
    The tolerances used for the fixed-point iteration and BiCGSTAB are set to $10^{-7}$ and $10^{-14}$, respectively.
  • $\bullet$
    Interpolation is performed by cubic Lagrange polynomials.
Figure 3
Figure 3: European put-on-the-average option price and the Greeks Delta and Gamma for the parameter set NIG0
displays the graphs of the option price function and its Greeks Delta and Gamma for the parameter set NIG0 from Table 1, where we have taken $N_{x}=400$. The Greeks have been approximated (at negligible computational cost) by applying the second-order central finite difference schemes described in Section 3.2. Table 3
Table 3: Numerical option prices for points $x$ near $\left(K,K\right)$
$\left(x^{\left(1\right)},x^{\left(2\right)}\right)$ VG0 VG1 NIG0 NIG1
$\left(90,90\right)$ 12.6534 10.1080 11.4067 11.5830
$\left(90,100\right)$ 10.6121 5.8462 7.8724 8.1529
$\left(90,110\right)$ 9.0136 3.0178 5.1023 5.4657
$\left(100,90\right)$ 10.4061 5.7637 7.8897 8.0910
$\left(100,100\right)$ 8.8015 2.9037 5.1186 5.3953
$\left(100,110\right)$ 7.5309 1.3893 3.1156 3.4311
$\left(110,90\right)$ 8.6181 2.8070 5.1393 5.3381
$\left(110,100\right)$ 7.3464 1.3062 3.1326 3.3735
$\left(110,110\right)$ 6.3290 0.6014 1.7937 2.0397
provides the numerical option prices for various points $x$ around $\left(K,K\right)$ and all four parameter sets from Table 1.
We next investigate the convergence behaviour of the numerical scheme. Let $\mathbf{x}_{N}$ denote the set of spatial grid points if $N_{x}=N$. For $x\in R_{x}$, let $\tilde{u}(x;N)$ denote the approximation of the exact solution value $u(x,T)$ obtained by the numerical scheme if $N_{x}=N$. More precisely, the vector $V^{N_{t}}$ generated by (3.27), (3.28) yields the approximation on the spatial grid $\mathbf{x}_{N}$ and cubic interpolation is employed whenever $x\notin\mathbf{x}_{N}$. We consider $\tilde{u}(x;N)$ with $N=400$ as the reference solution and study for $50\leq N\leq 200$ the total error defined by
$E\left(N\right)=\max\left\{|\tilde{u}(x;N)-\tilde{u}(x;400)|:x\in\mathbf{x}_{N}\textrm{\penalty 10000\ and\penalty 10000\ }x\in\left[0,3K\right]\times\left[0,3K\right]\right\}.$
Figure 4
Figure 4: Total error in $\left[0,3K\right]\times\left[0,3K\right]$
displays the total errors for all four parameter sets from Table 1. The quantity $p$ in the legend denotes the numerical order of convergence, which is computed by linear regression. Clearly, the favourable result is found that the numerical scheme achieves second-order convergence for each set of parameters.
5 Conclusions
In this paper, we have developed an effective numerical method for the valuation of European options under two-asset exponential Lévy models with particular attention to the infinite-activity case. Our method is based upon the ideas in Wang et al. (2007) for the one-asset case. A key part of our method is the tailored discretization of the non-local integral term, designed to handle singular measures under mild assumptions. The discretized integral term can subsequently be efficiently evaluated by FFT. For the discretization in time, the semi-Lagrangian Crank–Nicolson method is employed with a fixed-point iteration on the integral part. Numerical experiments for put-on-the-average options under Normal Tempered Stable processes indicate that our method achieves favourable second-order convergence whenever the exponential Lévy model has finite-variation.
A main topic for future research will be extending the proposed methodology to the valuation of American-style two-asset options under exponential Lévy models with infinite-activity, where the combination of the early-exercise feature and the non-local integral term poses additional challenges.
6 Acknowledgements
The authors acknowledge the support of the Research Foundation - Flanders (FWO) under grant G0B5623N and the FWO Scientific Research Network ModSimFIE (FWO WOG W001021N). The third author also acknowledges the financial support of the Research Foundation - Flanders (FWO) through FWO SAB K803124.
Appendix A $d$-dimensional Normal Tempered Stable process
The term $d$-dimensional Normal Tempered Stable process refers to a $d$-dimensional pure-jump compensated Lévy process $L$ with Lévy measure generated by subordinating a $d$-dimensional Brownian motion $B$ with a tempered stable subordinator $G$, i.e., a pure-jump process with almost surely non-decreasing trajectories. Such a process is defined by the following equation
$L\left(t\right)=B\left(G\left(t\right)\right)-\mathbb{E}\left[B\left(G\left(t\right)\right)\right]\qquad\text{with }L\left(0\right)=0.$
In our context, we will use this process to define the jump component of the logarithmic return in asset prices, i.e. we choose
$\int_{\mathbb{R}_{*}^{2}}z\tilde{\Pi}\left(dt,dz\right)=dL\left(t\right).$
A.1 Tempered Stable subordinator
A tempered stable subordinator is a non-compensated 1-sided tempered stable process $G$, which is characterized by the parameters $\delta,\lambda>0$ and $\alpha\in\left[0,1\right)$. For more details see Küchler & Tappe (2013). Table 4
Table 4: Main quantities of the Tempered Stable subordinator
Quantity Formula
Lévy measure $\ell_{G}\left(x\right)=\mathbb{I}_{x>0}\delta e^{-\lambda x}x^{-1-\alpha}$
Characteristic exponent $\psi_{G}\left(x\right)=\begin{cases}-\delta\ln\left(1-ix\lambda^{-1}\right)&\text{if }\alpha=0\\ \delta\Gamma\left(-\alpha\right)\left(\left(\lambda-ix\right)^{\alpha}-\lambda^{\alpha}\right)&\text{if }\alpha\in\left(0,1\right)\end{cases}$
Expected value $\mathbb{E}\left[G\left(1\right)\right]=\delta\frac{\Gamma\left(1-\alpha\right)}{\lambda^{1-\alpha}}$
Variance $\mathbb{V}\left[G\left(1\right)\right]=\delta\frac{\Gamma\left(2-\alpha\right)}{\lambda^{2-\alpha}}$
Density function $f_{G}\left(x\right)=\begin{cases}\mathbb{I}_{x>0}\frac{\lambda^{\delta}}{\Gamma\left(\delta\right)}x^{\delta-1}e^{-\lambda x}&\text{if }\alpha=0\\ \mathbb{I}_{x>0}\delta x^{-\frac{3}{2}}e^{-\left(\sqrt{\lambda}x-\delta\sqrt{\pi}\right)^{2}x^{-1}}&\text{if }\alpha=\frac{1}{2}\\ \text{not known analytically}&\text{else}\end{cases}$
shows the main quantities for such a process. Note that $G$ corresponds to the Gamma process for $\alpha=0$ and to the Inverse Gaussian process for $\alpha=\frac{1}{2}$.
A.2 Normal Tempered Stable process
Consider $B\left(t\right)=\eta t+\sqrt{\rho}\,W\left(t\right)$, where $W$ is a standard $d$-dimensional Wiener process, $\eta\in\mathbb{R}^{d}$ and $\sqrt{\rho}$ is the Cholesky decomposition of a given positive semi-definite symmetric matrix $\rho$, i.e., $\rho=\sqrt{\rho}\cdot\sqrt{\rho}^{\top}$. Adapting the results presented in Barndorff-Nielsen et al. (2001) and Rocha-Arteaga & Sato (2019, Chapter 4) [^8] The authors consider the more general case where the characteristic exponent of $L$ is defined as $\psi_{L}\left(\tau\right)=\int_{\mathbb{R}^{d}}\left(e^{i\tau^{\top}z}-1-i\tau^{\top}z\mathbb{I}_{\left\|z\right\|<1}\right)\ell_{L}\left(dz\right)$, while we consider the case where $\psi_{L}\left(\tau\right)=\int_{\mathbb{R}^{d}}\left(e^{i\tau^{\top}z}-1-i\tau^{\top}z\right)\ell_{L}\left(dz\right)$., we define a $d$-dimensional Normal Tempered Stable process as
$L\left(t\right)=B\left(G\left(t\right)\right)-ct,$ (A.1)
where $G$ is a Tempered Stable subordinator and $c=\mathbb{E}\left[B\left(G\left(t\right)\right)\right]=\delta\frac{\Gamma\left(1-\alpha\right)}{\lambda^{1-\alpha}}\eta$.
Table 5
Table 5: Main quantities of the Normal Tempered Stable process
Quantity Formula
Lévy measure $\ell_{L}\left(x\right)=\delta\Phi\left(x\mid\alpha,0\right)$
Characteristic exponent $\psi_{L}\left(x\right)=\begin{cases}-\delta\ln\left(\frac{\lambda-ix^{\top}\eta+\frac{1}{2}x^{\top}\rho x}{\lambda}\right)-ix^{\top}c&\text{if }\alpha=0\\ \delta\Gamma\left(-\alpha\right)\left(\left(\lambda-ix^{\top}\eta+\frac{1}{2}x^{\top}\rho x\right)^{\alpha}-\lambda^{\alpha}\right)-ix^{\top}c&\text{if }\alpha\in\left(0,1\right)\end{cases}$
Expected value $\mathbb{E}\left[L\left(1\right)\right]=0$
Variance $\mathbb{V}\left[L\left(1\right)\right]=\delta\frac{\Gamma\left(2-\alpha\right)}{\lambda^{2-\alpha}}\left(\rho\frac{\lambda}{1-\alpha}+\eta\eta^{\top}\right)$
Density function $f_{L}\left(x\right)=\begin{cases}\frac{\lambda^{\delta}}{\Gamma\left(\delta\right)}\Phi\left(x+c\mid-\delta,0\right)&\text{if }\alpha=0\\ \delta e^{2\delta\sqrt{\lambda\pi}}\Phi\left(x+c\mid\frac{1}{2},\delta^{2}\pi\right)&\text{if }\alpha=\frac{1}{2}\\ \text{not known analytically}&\text{else}\end{cases}$
shows the main quantities for such a process. Most of the formulae are expressed in terms of the function $\Phi$ which is given by
$\Phi\left(x\mid a,b\right)=2\sqrt{\frac{\left(\left\|\eta\right\|_{\rho}^{2}+2\lambda\right)^{a+\frac{d}{2}}}{\left(2\pi\right)^{d}\det\left[\rho\right]}}\frac{K_{a+\frac{d}{2}}\left(\sqrt{\left(\left\|\eta\right\|_{\rho}^{2}+2\lambda\right)\left(\left\|x\right\|_{\rho}^{2}+2b\right)}\right)}{\left(\sqrt{\left\|x\right\|_{\rho}^{2}+2b}\right)^{a+\frac{d}{2}}}e^{\left\langle\eta,x\right\rangle_{\rho}}$ (A.2)
where $K_{\nu}\left(\tau\right)=\frac{1}{2}\int_{0}^{\infty}y^{\nu-1}e^{-\frac{1}{2}\tau\left(y+y^{-1}\right)}dy$, for $\tau>0$, denotes the modified Bessel function of the second kind (see Schoutens (2003, Appendix A)), $\left\langle x,y\right\rangle_{\rho}=x^{\top}\rho^{-1}y$ and $\left\|x\right\|_{\rho}=\sqrt{\left\langle x,x\right\rangle}$ is its induced norm. We conclude this appendix with the following proposition.
Proposition A.1.
Consider a Lévy measure $\ell$ over $\mathbb{R}^{d}_{*}=\mathbb{R}^{d}\setminus\{0\}$. Assume that there exist constants $A_{\ell}$ and $B_{\ell}$, and for any given $h>0$ a constant $C_{\ell}(h)$ such that
$\begin{cases}\ell\left(z\right)\leq C_{\ell}\left(h\right)\left\|z\right\|_{\rho}^{-A_{\ell}-d}&\text{for any }z\in\mathbb{R}^{d}_{*}\text{ such that }\left\|z\right\|_{\rho}\in\left(0,h\right]\\ \ell\left(z\right)=O\left(e^{-B_{\ell}\left\|z\right\|_{\rho}}\right)&\text{as }\left\|z\right\|_{\rho}\rightarrow\infty.\end{cases}$
Then, for a Normal Tempered Stable process these constants are given by
$A_{\ell}$ $=2\alpha$
$B_{\ell}$ $=\sqrt{\left\|\eta\right\|_{\rho}^{2}+2\lambda}-\left\|\eta\right\|_{\rho}$
$C_{\ell}\left(h\right)$ $=\frac{2^{\alpha}\delta\Gamma\left(\alpha+\frac{d}{2}\right)}{\sqrt{\pi^{d}\det\left[\rho\right]}}e^{h\left\|\eta\right\|_{\rho}}.$
Proof.
Denote $c_{1}=\sqrt{\left\|\eta\right\|_{\rho}^{2}+2\lambda}$ and $c_{2}=2\delta c_{1}^{\alpha+\frac{d}{2}}\left(2\pi\right)^{-\frac{d}{2}}\det\left[\rho\right]^{-\frac{1}{2}}$. The function $\overline{\ell}:\mathbb{R}_{>0}\rightarrow\mathbb{R}$, defined by
$\overline{\ell}\left(\tau\right)=c_{2}\tau^{-\alpha-\frac{d}{2}}K_{\alpha+\frac{d}{2}}\left(c_{1}\tau\right)e^{\tau\left\|\eta\right\|_{\rho}},$
represents a radial upper bound for $\ell$, since by the Cauchy-Schwarz inequality $\ell\left(z\right)\leq\overline{\ell}\left(\left\|z\right\|_{\rho}\right)$ for any $z\in\mathbb{R}^{d}_{*}$. Applying the following well known inequality for the modified Bessel function of the second kind
$\tau^{\nu}K_{\nu}\left(\tau\right)\leq\Gamma\left(\nu\right)2^{\nu-1}\qquad\text{for any }\tau,\nu>0,$
we readily get
$\overline{\ell}\left(\tau\right)\leq c_{2}c_{1}^{-\alpha-\frac{d}{2}}2^{\alpha+\frac{d}{2}-1}e^{\tau\left\|\eta\right\|_{\rho}}\Gamma\left(\alpha+\frac{d}{2}\right)\tau^{-2\alpha-d}.$
For $\tau=\left\|z\right\|_{\rho}$ with $\left\|z\right\|_{\rho}\in\left(0,h\right]$ we easily deduce the stated expressions for $A_{\ell}$ and $C_{\ell}(h)$. Invoking the following asymptotic behaviour of the modified Bessel function of the second kind
$K_{\nu}\left(\tau\right)=O\left(\tau^{-\frac{1}{2}}e^{-\tau}\right)\qquad\text{as }\tau\rightarrow\infty,$
it follows that
$\overline{\ell}\left(\tau\right)=O\left(\tau^{-\alpha-\frac{d}{2}-\frac{1}{2}}e^{-\tau\left(c_{1}-\left\|\eta\right\|_{\rho}\right)}\right).$
Since $\tau^{-\alpha-\frac{d}{2}-\frac{1}{2}}=O\left(1\right)$ as $\tau\rightarrow\infty$, we obtain the expression for $B_{\ell}$. ∎
Appendix B Summation operator as a circulant matrix-vector multiplication
Let $N_{z}=1$, $N_{y}^{-}=0$ and $N_{y}^{+}=1$. Then $\sharp\text{out}=N_{y}^{+}+N_{y}^{-}+1=2$ and $\sharp\text{in}=2N_{z}+N_{y}^{+}+N_{y}^{-}=3$. The quadrature matrix $\Omega$, whose entries are the coefficients $\omega$ defined in (3.5), is given by
$\Omega=\left[\begin{array}[]{cc}\omega_{-1,-1}&\omega_{-1,0}\\ \omega_{0,-1}&\omega_{0,0}\end{array}\right]\in\mathbb{R}^{2N_{z}\times 2N_{z}}.$
The first row of the circulant matrix $C$ is defined according to
$C_{1,\cdot}=\text{vec}\left(\left[\begin{array}[]{ccc}\omega_{-1,-1}&\omega_{-1,0}&0\\ \omega_{0,-1}&\omega_{0,0}&0\\ 0&0&0\end{array}\right]\right)\in\mathbb{R}^{\left(\sharp\text{in}\right)^{2}\times 1},$
while the entire matrix is
$C=\left[\begin{array}[]{ccccccccc}{\color[rgb]{1,0,0}\omega_{-1,-1}}&{\color[rgb]{1,0,0}\omega_{0,-1}}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}\omega_{-1,0}}&{\color[rgb]{1,0,0}\omega_{0,0}}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}\omega_{-1,-1}}&{\color[rgb]{1,0,0}\omega_{0,-1}}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}\omega_{-1,0}}&{\color[rgb]{1,0,0}\omega_{0,0}}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}\\ 0&0&\omega_{-1,-1}&\omega_{0,-1}&0&\omega_{-1,0}&\omega_{0,0}&0&0\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}\omega_{-1,-1}}&{\color[rgb]{1,0,0}\omega_{0,-1}}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}\omega_{-1,0}}&{\color[rgb]{1,0,0}\omega_{0,0}}&{\color[rgb]{1,0,0}0}\\ {\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}\omega_{-1,-1}}&{\color[rgb]{1,0,0}\omega_{0,-1}}&{\color[rgb]{1,0,0}0}&{\color[rgb]{1,0,0}\omega_{-1,0}}&{\color[rgb]{1,0,0}\omega_{0,0}}\\ \omega_{0,0}&0&0&0&0&\omega_{-1,-1}&\omega_{0,-1}&0&\omega_{-1,0}\\ \omega_{-1,0}&\omega_{0,0}&0&0&0&0&\omega_{-1,-1}&\omega_{0,-1}&0\\ 0&\omega_{-1,0}&\omega_{0,0}&0&0&0&0&\omega_{-1,-1}&\omega_{0,-1}\\ \omega_{0,-1}&0&\omega_{-1,0}&\omega_{0,0}&0&0&0&0&\omega_{-1,-1}\end{array}\right]\in\mathbb{R}^{\left(\sharp\text{in}\right)^{2}\times\left(\sharp\text{in}\right)^{2}}.$
The entries in the first, second, fourth, and fifth rows (highlighted in red) correspond to the matrix $\tilde{I}C\in\mathbb{R}^{\left(\sharp\text{out}\right)^{2}\times\left(\sharp\text{in}\right)^{2}}$.